Sara and Josh were playing a board game in which they rolled two dice.
Rolling doubles (that is, rolling the same number on both sice) lets take
an extra turn. Josh roleed two 3s and then two 5s. As she was getting ready
to take another extra turn, Sara said 'The chances of you getting doubles
again are next to nothing' Is craig correct? Explain your answer

Answer by venugopalramana(3286)

Sara and Josh were playing a board game in which they rolled two dice.
Rolling doubles (that is, rolling the same number on both sice) lets take
an extra turn. Josh roleed two 3s and then two 5s. As she was getting ready
to take another extra turn, Sara said 'The chances of you getting doubles
again are next to nothing' Is craig correct? Explain your answer
PROBABILITY OF GETTING A DOUBLE:
IN I THROW OF 2 DICE..
TOTAL POSSIBILITUES =6*6=36
POSSIBILITIES OF GETTING DOUBLE=6
P1=PROBABILITY=6/36=1/6
IN II THROW 2 DICE
THESE ARE INDEPENDENT EVENTS.HENCE AGAIN
P2=1/6
SIMILARLY AT ANT TRIAL
P=1/6...
HENCE IF WE SPECIFY IN THE BEGINING GETTING 3 DOUBLES IN 3 THROWS THEN
P=(1/6)*(1/6)*(1/6)=1/216 WHICH IS LOW.
BUT THE POINT IS THAT HAVING ALREADY ACHIEVED 2 DOUBLES OR NOT FOR THAT MATTER,EACH THROW IS INDEPENDENT AND THE PROBABILITY IN THAT PARTICULAR THROW ONLY IS NEITHER MORE NOR LESS...SO PREDICTING THAT THE PROBABILITY OF GETTING A DOUBLE AGAIN REPEAT AGAIN IS NEXT TO NOTHING IS NOT REPEAT NOT CORRECT.IT IS NOT AFFECTED BY PREVIOUS RESULTS.ONLY WHEN YOU TRY TO PREDICT IN THE BEGINING THAT ACHIEVING 3 DOUBLES IN 3 THROWS IS VERY SMALL IS OK